# Are You Getting What You Pay For? Auditing Model Substitution in LLM APIs

### v1

@@ 4.3 Hardware-assisted verification: Trusted Execution Environment (TEE)
Method.
Trusted Execution Environment (TEE) technology provides a secure and isolated zone within the processor, safeguarding data processed within the secure enclave from unauthorized access and alterations. This allows users to gain a clear understanding of how their data is handled and processed, as the TEE guarantees integrity and confidentiality throughout computations by hardware. NVIDIA Confidential Computing (NVIDIA, 2023) brings TEE technology to their latest GPUs in the Hopper and Blackwell architectures, making model inference within TEEs possible with minimal performance cost. This method enables end users to verify the authenticity of the model in use, ensuring that the advertised model is indeed operational as the backend model. In addition to ensuring the integrity of the model, this technology can also provide privacy guarantees for user input with proper settings. Appendix C includes more details about building an LLM inference API endpoint with TEE. This method shifts trust assumptions from the API provider’s operational honesty to the integrity of the hardware manufacturer.
Evaluation.
Metric Concurrent requests # TEE no TEE Overhead
First token latency (ms) 1 71 65 9.23%
@@
Overall throughput (token/s) 1 99 117 15.38%
@@
Table 4: Performance comparison.
To evaluate the performance overhead of using TEEs, we conduct a performance benchmark for a vLLM API endpoint with and without TEE. We host a Meta-Llama-3-8B-Instruct model with single H100 GPU and enable TLS for the HTTP endpoint. As can be observed in Table 4, the overheads are relatively low, especially the overall throughput under concurrent requests, which has an overhead of less than 3%. In real-world scenarios, API providers serve multiple users simultaneously and receive a large number of concurrent requests. They charge users based on the number of tokens. Therefore, the overall throughput under concurrent requests is the metric that API providers care about most.
5 Discussion and Conclusion
Our investigation into auditing model substitution in LLM APIs reveals a challenging landscape for users seeking verification.
Limitations of software-based verification.
Methods relying solely on text output (classification, identity prompting, MMD, benchmarks) face significant hurdles. They struggle to detect subtle substitutions like quantization (Table 2, Table 3) and can be undermined by adversarial strategies. Randomized substitution significantly degrades the power of statistical tests like MMD (Figure 1), while benchmark evasion can trivially defeat methods relying on known prompts (identity prompting, benchmarks). Furthermore, the effectiveness of benchmarks is highly sensitive to undisclosed decoding parameters like temperature (Figure 2), making reliable comparisons difficult without full transparency.
Strength and weakness of log probability analysis.
Access to token log probabilities provides a much stronger signal (Figure 3). Logprob distributions are highly sensitive to model weights, allowing reliable detection of even quantization or minor version changes, assuming the auditor has a reference $M_{spec}$ instance. However, this method’s applicability is entirely dependent on the provider offering logprobs access (Table 1). Providers can easily thwart this technique by simply not providing logprobs (limiting information disclosure) or potentially serving correct logprobs only for detected audit queries (benchmark evasion). In addition, optimization techniques like continuous batching may affect the results and the need for the user to perform reference computations also adds overhead.
The promise and challenge of TEEs.
Hardware-assisted verification using TEEs (Section 4.3) offers the most robust solution against software-level substitution and evasion tactics. It provides a strong integrity guarantee about the code and data (model weights) being executed. Performance overhead appears manageable, especially under high concurrency (Table 4). However, the primary barrier is infrastructure dependency and provider adoption. TEEs require specific hardware and software setup, and providers must be willing to offer attestation services. Establishing trust also requires verifiable measurements for the inference stack and model weights, necessitating open-source software components and mechanisms for distributing or verifying model hashes.
Overall landscape and future directions.
Currently, no universally applicable and robust software-based method exists for users to reliably audit model substitution against a determined adversary. Log probability verification is the strongest software technique but it is not sufficiently accurate and relies on API features. TEEs offer a technical path to provable integrity but require ecosystem support. This highlights a critical need for:
  • Increased transparency: Providers could voluntarily increase transparency by consistently providing detailed metadata like log probabilities, exact model versions (including quantization details), and potentially TEE attestations.
  • More robust auditing techniques: Research into statistical methods less susceptible to randomization or requiring fewer samples, potentially combining multiple weak signals, could be valuable. Techniques analyzing finer-grained output properties beyond standard text features might also yield improvements.
  • Standardized protocols: Development of standardized protocols for verifiable LLM inference, perhaps incorporating lightweight cryptographic commitments or attested execution, could build trust between users and providers.
  • User awareness and tools: Empowering users with tools and knowledge about potential substitutions and available verification methods is crucial.
@@ Appendix B Model stealing and embedding verification
Background.
Researchers have explored extracting model information (e.g., weights, architecture details) from black-box APIs (Carlini et al., 2024; Finlayson et al., 2024). While successful extraction could reveal substitutions, these methods often require a vast number of queries, significant computational resources, and may still yield incomplete information, making them impractical for routine auditing by typical users. Finlayson et al. (2024) specifically demonstrated how logits analysis could expose model details, which we consider as a potential verification method for detecting model substitutions.
Method.
The embedding size detection relies on the fact that logits generated by an LLM are restricted to a $d$-dimensional subspace of the full $v$-dimensional output space where $d$ is the hidden embedding size. So the auditor can analyze a set of logit outputs from different prompts and reconstruct this subspace and the hidden embedding size.
Given a set of $n$ output logit vectors $\{\ell_{i}\}_{i=1}^{n}$ (or log probabilities, since it preserves linear relationship up to a constant addition), we construct a logit matrix $L$ with each vectors as its columns. Applying SVD to $L$ yields $L=U\Sigma V^{T},$ where $U\in\mathbb{R}^{v\times v}$, $\Sigma\in\mathbb{R}^{v\times n}$ (with singular values $\lambda_{1}\geq\lambda_{2}\geq\dots$), and $V\in\mathbb{R}^{n\times n}$. The magnitude of the singular values will drastically decreases after the first $d$ dimensions. Formally, the embedding size $d$ can be identified by detecting the singular value index at which the magnitude drops the most: $d=\arg\max_{i}(\log\lambda_{i}-\log\lambda_{i+1}),$ Once identified, the $d$-dimensional subspace is constructed using the first $d$ left singular vectors from $U$: $U_{d}=[u_{1},u_{2},\dots,u_{d}]\in\mathbb{R}^{v\times d},$ forming a unique signature of the model. Empirical evaluations confirm that this method accurately recovers the embedding dimensions. Furthermore, monitoring this subspace allows auditors to detect subtle changes, such as hidden system prompts, fine-tuning, or entire model substitutions from the service provider.
Weakness.
This verification method is impractical because no well-known service provider gives full log probability access.

### v2

@@ 4.3 Hardware-assisted verification: Trusted Execution Environment (TEE)
Method. Trusted Execution Environments (TEEs), such as those provided by NVIDIA’s Confidential Computing on Hopper and Blackwell GPUs (NVIDIA, 2023), offer a paradigm shift. A TEE creates a hardware-isolated enclave where both the model weights and the inference code are protected from the host system. The TEE can produce a cryptographic attestation report that includes measurements (hashes) of the loaded model and execution code. An end-user can cryptographically verify this report to gain absolute certainty that the intended model is running unmodified within the specified, trusted software environment.
Evaluation. The primary concern with TEEs is performance overhead. We benchmark a vLLM inference endpoint for Llama-3-8B running on a single H100 GPU, both with and without a TEE enabled. We measure first-token latency and overall throughput under single-request and high-concurrency (64 requests) workloads.
Metric
Concurrent
requests #
TEE no TEE Overhead
First token latency
1 71 65 9.23%
@@
Overall throughput
(token/s)
1 99 117 15.38%
@@
Table 4: Performance comparison for a vLLM endpoint with/without TEE (Meta-Llama-3-8B-Instruct, single H100, TLS enabled).
Analysis. Unlike software methods, TEEs are not vulnerable to computational attacks, as trust is anchored in hardware. The security guarantee is cryptographic, not statistical. Our evaluation in Table 4 shows that this strong guarantee comes at a modest and practical cost. The overhead for first-token latency was 9-16%, while the drop in overall throughput under a high-concurrency load was only 2.88%. This demonstrates that TEEs offer a deployable, high-assurance solution to the model substitution problem, effectively closing the verification gap left by software-only techniques.
5 Discussion
Our systematic evaluation reveals a clear hierarchy of verification capabilities:
Text-only methods are insufficient. Methods that only look at output distributions (classifiers, identity prompts, MMD, benchmarks) either need many queries, get confounded by benign variance, or can be evaded with some level of quantized model substitution. These approaches can give a rough signal but do not provide strong integrity guarantees. More generally, limited information access forces auditors to rely on large sample sizes to identify weak statistical patterns.
Metadata-based methods are fragile. Logprobs offer more sensitivity: in principle, per-token likelihoods can reveal discrepancies with far fewer samples than text-only outputs. However, in practice, backend nondeterminism destroys this advantage: variance across frameworks, batching, and hardware is often larger than the effect of quantization or fine-tuning. TopLoc improves robustness by producing compact activation-based proofs that are less sensitive to such variability, but it has no guarantees under significantly different inference settings, and still requires local recomputation of $M_{\text{spec}}$, making it infeasible for proprietary deployments. In practice, both logprobs and TopLoc collapse without access to the reference model.
Hardware-backed attestation is strongest. Trusted Execution Environments (TEEs) avoid these weaknesses: they guarantee integrity without exposing internals or requiring the auditor to hold the weights. Overheads are modest (Table 4), and attestation cryptographically binds the stack and model hash to the outputs. Adoption remains limited by operational complexity and weak provider incentives, but TEEs are currently the best compromise between verifiability, efficiency, and protection of proprietary models.
Actionable recommendations: (1) For users: Request attestation proofs when available; for critical applications, prefer providers offering TEE-backed services. (2) For API providers: Consider TEE deployment for premium tiers; increase transparency through model versioning and metadata disclosure. (3) For policymakers: Develop standards for API transparency and consider requiring attestation for regulated applications.
6 Conclusion
The model substitution problem represents a fundamental challenge in the current LLM ecosystem, where economic incentives misalign with user trust. Our work provides both a comprehensive analysis of why existing approaches fail and a concrete path forward through hardware attestation. As LLM APIs become increasingly critical infrastructure, ensuring computational integrity through TEEs may become as essential as HTTPS was for web security.
Ethics Statement
This study focuses on auditing techniques for detecting model substitution in LLM APIs. All experiments are conducted using public datasets and benchmarks without involving human subjects or personal data, and API queries are issued within the provider’s terms of service and rate limits. The attacks we discuss against text-output–based verification are not presented to encourage misuse, but to motivate the research community to develop stronger auditing methods under the same level of API access. Our goal is to improve transparency and trust in LLM services by highlighting weaknesses in current auditing approaches.
@@ Appendix B Model stealing and embedding verification
Background. Researchers have explored extracting model information (e.g., weights, architecture details) from black-box APIs (Carlini et al., 2024; Finlayson et al., 2024). While successful extraction could reveal substitutions, these methods often require a vast number of queries, significant computational resources, and may still yield incomplete information, making them impractical for routine auditing by typical users. Finlayson et al. (2024) specifically demonstrate how logits analysis could expose model details, which we consider as a potential verification method for detecting model substitutions.
Method. The embedding size detection relies on the fact that logits generated by an LLM are restricted to a $d$-dimensional subspace of the full $v$-dimensional output space where $d$ is the hidden embedding size. So the auditor can analyze a set of logit outputs from different prompts and reconstruct this subspace and the hidden embedding size.
Given a set of $n$ output logit vectors $\{\ell_{i}\}_{i=1}^{n}$ (or log probabilities, since it preserves linear relationship up to a constant addition), we construct a logit matrix $L$ with each vector as its columns. Applying SVD to $L$ yields $L=U\Sigma V^{T},$ where $U\in\mathbb{R}^{v\times v}$, $\Sigma\in\mathbb{R}^{v\times n}$ (with singular values $\lambda_{1}\geq\lambda_{2}\geq\dots$), and $V\in\mathbb{R}^{n\times n}$. The magnitude of the singular values will drastically decreases after the first $d$ dimensions. Formally, the embedding size $d$ can be identified by detecting the singular value index at which the magnitude drops the most: $d=\arg\max_{i}(\log\lambda_{i}-\log\lambda_{i+1}),$ Once identified, the $d$-dimensional subspace is constructed using the first $d$ left singular vectors from $U$: $U_{d}=[u_{1},u_{2},\dots,u_{d}]\in\mathbb{R}^{v\times d},$ forming a unique signature of the model. Empirical evaluations confirm that this method accurately recovers the embedding dimensions. Furthermore, monitoring this subspace allows auditors to detect subtle changes, such as hidden system prompts, fine-tuning, or entire model substitutions from the service provider.
Weakness. This verification method is impractical because no well-known service provider gives full log probability access.